Weatherwise is a magazine published by the American Meteorological Society. One issue gives a rating system used to classify Nor'easter storms that frequently hit New England and can cause much damage near the ocean. A severe storm has an average peak wave height of μ = 16.4 feet for waves hitting the shore. Suppose that a Nor'easter is in progress at the severe storm class rating. Peak wave heights are usually measured from land (using binoculars) off fixed cement piers. Suppose that a reading of 34 waves showed an average wave height of x = 17.5 feet. Previous studies of severe storms indicate that σ = 3.5 feet. Does this information suggest that the storm is (perhaps temporarily) increasing above the severe rating? Use α = 0.01.

(a) What is the level of significance? ____________________________

(b) What is the value of the sample test statistic? (Round your answer to two decimal places.) ______________

(c) Find the P-value. (Round your answer to four decimal places.) ___________________
Answer #1

Solution :

This is the right tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 16.4

Ha : \mu > 16.4

\bar x = 17.5

\mu = 16.4

\sigma = 3.5

n = 34

(a) level of significance = 0.01

(b)

Test statistic = z

= (\bar x - \mu) / \sigma / \sqrtn

= (17.5 - 16.4) /3.5 / \sqrt34

= 1.83

Test statistic = 1.83

(c)

P(z > 1.83) = 1 - P(z < 1.83) = 1 - 0.9664 = 0.0336

P-value = 0.0336

\alpha = 0.01

P-value > \alpha

Fail to reject the null hypothesis .

There is no evidence to suggest that the storm is (perhaps temporarily) increasing above the severe rating .